# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

### v3

@@ 5.3 Ablation Study
Table 6: Fixed vs. Dynamic Bandit Variants over GPT-4o-mini. #US: Number of Unique Strategies.
Figure 2: We normalize the inference costs (based on token usage) in a range from 0 to 1 for easier comparison. We use gpt-4o-mini as the backend model for all the methods in this figure. For each method, key hyper-parameters (e.g., ${N}$ in Best of N, or tree size in ToT) are tuned to yield a baseline (lower point) and an extended (upper point) configuration, with dashed lines connecting these bounds.
In this section, we conduct an ablation study to analyze each component contribution of Meta-Reasoner. In specific, we consider the following setup: (1) w/o progress report: we replace the progress reporting process with directly considering the entire CoT history without summarization; (2) w/o MAB: instead of using MAB to select the proper strategy, we directly leverage an LLM to the decision making to provide the proper strategy for LRM reasoning. In Table 5, we show that when removing progress reporting (“w/o Progress Report”), the overall performance moderately degrades and we hypothesize it is due to the concise intermediate summarizations can help the Meta-reasoner only consider the high-level strategy instead of being confused with too much details of the reasoning process. We also find that removing the MAB brings a more pronounced effect, especially when strategy selection falls back to a direct chain-of-thought approach (“w/o MAB (CoT)”). It verifies the effect of our meta-reasoner module to help the model stay on track for getting an optimal solution. In Table 6, we compare fixed and dynamic bandit variants on the game of 24 and theoremQA. We find that using a fixed set of strategies (e.g., $K=3$ and $K=5$) yields lower performance compared to the dynamic approach which adaptively explores more strategies (shown by larger unique strategies). The results highlight the benefit of flexibly allocating diverse reasoning strategies using LLM in-context learning capabilities.
5.4 Analysis
Figure 3: Cumulative reward of different settings across iteration. We compare our method using LinUCB with baseline (direct arm selection), and random search methods across two tasks—Game of 24 (top row) and TheoremQA (bottom row) using GPT-40-mini (left) and Gemini-Exp-1206 (right).
Inference Efficiency.
A common way to measure computational efficiency is wall-clock time, as it directly reflects real-world latency during inference. However, in our API-based setup, wall-clock time is unreliable due to external factors such as network latency and server load, which introduce noise unrelated to the model’s intrinsic computational cost. Therefore, we use token usage as a more stable and meaningful proxy, since it directly correlates with the amount of computation performed and is widely adopted in recent works Liu et al. (2025); Li et al. (2025b). The detailed instruction can be found in Appendix B.1.
Figure 2 presents normalized inference costs (measured by token usage) alongside accuracy for various models and prompting strategies. Basic models like GPT-4o-mini and GPT-4o exhibit low accuracy and minimal cost, whereas advanced methods such as ToT Yao et al. (2023) and MACM Lei et al. (2024) improve accuracy at the expense of increased token consumption. Our method achieves a strong balance, achieving high accuracy with moderate token usage, outperforming MACM, which is more costly yet less accurate. Proprietary models like o1-mini and o1-preview attain slightly higher accuracy but incur the highest cost, underscoring their greater resource demands. Overall, Meta-Reasoner provides a cost-effective and scalable solution for reasoning-intensive tasks.
Effectiveness of Meta-reasoner.
Figure 3 demonstrates the cumulative rewards across iterations. We compare our MAB-based approach with a baseline that directly prompts an LLM to select an arm (or “strategy”), referred to as Baseline (Direct Arm Selection); the prompt details are in Figures 47. Results show that the MAB-based meta-reasoner (using LinUCB Li et al. (2012)) consistently outperforms both direct LLM decision-making and random search across two tasks (Game of 24 and TheoremQA) and two model scales (GPT-4o-mini and Gemini-Exp-1206). While direct LLM prompting yields reasonable initial performance and random search requires minimal setup, neither approach effectively balances exploration and exploitation. In contrast, the MAB updating strategy leverages feedback from prior iterations to adaptively refine action selection (e.g., choosing an appropriate strategy based on CoT reasoning), steadily increasing cumulative rewards.
6 Conclusion
@@
Limitations
Our proposed Meta-Reasoner framework, while effective at enhancing inference-time reasoning, remains limited to text-based problems and struggles to address tasks requiring other modalities, such as geometry. Overcoming these challenges calls for further advancements in the model’s cognitive capabilities.
Appendix A Baselines
We consider several established prompting methods as baselines as follows:
  • Chain-of-thought (CoT) Wei et al. (2022): A prompting technique that encourages models to generate intermediate reasoning steps to enhance problem-solving capabilities.
  • Self-Consistent Chain of Thought (SC-CoT) Wang et al. (2022): An extension of CoT that improves reasoning consistency by generating multiple reasoning chains and selecting the most consistent answer.
  • Multi-Chain Reasoning (MCR) Yoran et al. (2024): enhances SC-CoT by having another LLM to assess and integrate content among the sampled reasoning chains to generate the final consistent answer.
  • Tree of Thoughts (ToT) Yao et al. (2023): A method that explores multiple reasoning paths in a tree structure, allowing the model to consider various possibilities before arriving at a conclusion by tree search algorithms.
  • Reflexion Shinn et al. (2024): A framework that enables models to reflect on their reasoning process, iteratively refining their answers based on feedback.
  • MACM Lei et al. (2024): A multi-agent system to refine the reasoning based on iterative condition mining.
Appendix B Computational Efficiency
B.1 Why Measure Efficiency Through Token Usage?
Evaluating the computational efficiency of LLMs is a critical yet nuanced task. A natural metric is wall-clock time, which directly reflects the real-world latency experienced during LLM inference. However, when interacting with LLMs primarily through API calls, as what we do in this paper, wall-clock time measurements become unreliable and difficult to interpret. This unreliability arises from several uncontrollable factors, including network latency, server-side load balancing, and dynamic optimizations on the provider’s infrastructure. These external variables introduce noise and variability that obscure the intrinsic computational characteristics of the reasoning method itself.
To address these challenges, we adopt token usage as a more stable and conceptually meaningful proxy for computational cost. Token consumption directly correlates with the amount of computation performed by the LLM and is widely recognized in recent literature as a standard efficiency metric Liu et al. (2025); Li et al. (2025b). By focusing on token reduction, we provide a consistent basis for comparing reasoning strategies that is less susceptible to external system variability.
B.2 Iterative Reasoning and Memory Management
Unlike traditional CoT methods that generate reasoning process in a single pass, Meta-Reasoner pauses generation periodically to evaluate the reasoning progress and adaptively switch the reasoning strategies using a multi-armed bandit-based algorithm. This iterative design adds additional memory overhead due to maintaining context across multiple LLM calls. To mitigate this, we summarize the entire reasoning history into compact progress reports (§4.2), greatly reducing the token load per iteration. For example, a typical 1000-token CoT step may consume approximately 500MB of KV cache (e.g., using LLaMa-2-7B in float16 precision), while our method’s 5 iteration with 100-token steps peak at roughly 50MB each. Additionally, early pruning of unlikely reasoning paths further controls memory usage.
B.3 Managing Reward Model Overhead
To minimize the computational cost of the reward model, our design relies on summarized progress reports rather than the full reasoning history, thereby reducing token usage and associated processing overhead. The reward computation is further optimized by employing the LinUCB algorithm, a MAB method whose complexity scales linearly with the number of strategies. Since the number of strategies remains relatively small—typically between 3 and 5 in fixed settings and up to 10 to 15 in dynamic settings—the computational burden remains manageable. As demonstrated in Figure 2, the savings achieved by avoiding unproductive reasoning paths significantly outweigh the modest overhead introduced by the strategy selection process, resulting in an overall efficient reward computation framework.
Appendix C Stability of Dynamic Strategy Generation
Dynamically adding or refining strategies, as discussed in §4.3 (dynamic contextual bandit), enhances flexibility but risks instability. To address this trade-off, our approach incorporates several key mechanisms:
We begin with a small set of verified strategies (Table 1) to establish a stable foundation and prevent arbitrary generation. Building on this, the LLM generate contextually relevant and precise strategies, like “Pause to clarify and disambiguate the current reasoning” and “Simplify by decomposing the task into smaller, manageable sub-tasks”, demonstrate how the meta-reasoner effectively targets specific reasoning bottlenecks.
Second, we employ a contextual multi-armed bandit to balance exploration and exploitation, filtering out suboptimal strategies during LLM inference time. As shown in Table 6, the dynamic bandit variant achieves 89% accuracy on task Game-of-24, significantly surpassing fixed strategy sets (65%–72%). This highlights the dynamic bandit’s ability to prioritize effective strategies while discarding less useful ones.
Third, the reward function described in §4.3 provides immediate feedback on strategy performance, enabling rapid deprioritization of confusing or unproductive strategies. Figure 3 shows a consistent increase in cumulative rewards over iterations, indicating that the system maintains coherent reasoning as it incorporates new strategies.
Together, these mechanisms enable the dynamic contextual bandit to maintain stability without compromising adaptability. The observed performance gains on both the Game-of-24 and TheoremQA datasets (Table 6) show that the benefits of dynamic strategy generation substantially outweigh potential instability concerns in practice.
Algorithm 1 Meta-Reasoner: Meta-Reasoning with Contextual Multi-Armed Bandits
0:  LRM $M$, bandit $\mathcal{B}$, initial strategy set $\mathcal{A}_{1}$, maximum rounds $T$
0:  Final answer $A_{\mathrm{final}}$
1:  $C_{0}\leftarrow\emptyset$;  .$\mathcal{B}.\mathrm{Initialize}(\mathcal{A}_{1})$
2:  $G_{0}\leftarrow$ default strategy
3:  for $t=1$ to $T$ do
4:     if $t>1$ then
5:        $P_{t-1}\leftarrow f(C_{t-1})$ // Summarize the existing CoT
6:        $x_{t-1}\leftarrow\mathrm{FeatureExtract}(P_{t-1})$ // Extract features for context
7:        (Optional): $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\{\text{new strategies}\}$ // Update strategy set dynamically
8:        $a_{t-1}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}_{\mathcal{B}}(x_{t-1},a)$ // Select strategy using bandit
9:        $G_{t}\leftarrow a_{t-1}$ // Set current guidance
10:     else
11:        $G_{t}\leftarrow G_{0}$ // Use default guidance for the first iteration
12:     end if
13:     $s_{t}\leftarrow M(C_{t-1},G_{t})$ // Generate new CoT with integrated guidance
14:     $C_{t}\leftarrow C_{t-1}\cup\{s_{t}\}$ // Append new reasoning step to the CoT
15:     $r_{t}\leftarrow\mathrm{ComputeReward}(C_{t})$ // Compute reward based on the updated CoT
16:     if $t>1$ then
17:        .$\mathcal{B}.\mathrm{Update}(x_{t-1},a_{t-1},r_{t})$ // Update bandit with observed feedback
18:     end if
19:     if termination condition met then
20:        break
21:     end if
22:  end for
23:  $A_{\mathrm{final}}\leftarrow\mathrm{ExtractAnswer}(C_{t})$
24:  return  $A_{\mathrm{final}}$
Figure 4: Prompt Demonstration (Page-1)
Figure 5: Prompt Demonstration (Page-2)
Figure 6: Prompt Demonstration (Page-3)
Figure 7: Prompt Demonstration (Page-4)

### v4

@@ 5.4. Analysis
5.4.2. Inference Efficiency.
We measure the inference efficieny of our proposed method. In Figure 4, we calculate the average inference time on different tasks across different models and methods. It demonstrates Meta-Reasoner’s significant advantages over existing reasoning frameworks across multiple benchmarks. Compared to o1-preview in the zero-shot setting, our method achieved roughly 51-55% inference time reduction while maintaining comparable performance. Among reasoning methods using identical base models, Meta-Reasoner consistently outperformed alternatives: requiring 29-33% fewer inference time than ToT (Yao et al., 2023), Reflexion (Shinn et al., 2023), MACM (Lei et al., 2024) and Best-of-N techniques.We further provide the demonstration of the token usage comparison in Figure 5. The efficiency results positions Meta-Reasoner as a scalable solution for complex reasoning tasks where computational resources are constrained.
(a) Token Usage on Game of 24 Task.
(b) Token Usage on TheoremQA Task
Figure 5. Token Usage Across Different Methods on Game-of-24 (Yao et al., 2023) and TheoremQA (Chen et al., 2023) Tasks.
5.4.3. Iterative Reasoning and Memory Management.
Unlike traditional CoT methods that generate reasoning process in a single pass, Meta-Reasoner pauses generation periodically to evaluate the reasoning progress and adaptively switch the reasoning strategies using a multi-armed bandit-based algorithm. This iterative design adds additional memory overhead due to maintaining context across multiple LLM calls. To mitigate this, we summarize the entire reasoning history into compact progress reports (§4.2), greatly reducing the token load per iteration. As demonstrated in Figure 4, the savings achieved by avoiding unproductive reasoning paths significantly outweigh the modest overhead introduced by the strategy selection process, resulting in an overall efficient reward computation framework.
5.4.4. Generalizability of Meta-Reasoner.
To analyze the generalizability of our proposed method, we conduct additional experiment on creative writing, a domain that is quite different from math or logical reasoning. Following the instructions from ToT (Yao et al., 2023), we test our method on creative writing tasks where the input consists of four random sentences, and the output is a passage of four paragraphs ending with those sentences. We used the same evaluation metrics as ToT, which involves a GPT-4 zero-shot prompt providing a 1-10 scalar coherency score.
Table 7. Performance on Creative Writing Task (Yao et al., 2023).
Method GPT-4 Coherency Score
IO (Input-Output) 6.19
IO + Prompt Refined (k=5) 7.67
CoT (Chain of Thought) (Wei et al., 2022) 6.93
ToT (Tree of Thoughts) (Yao et al., 2023) 7.56
ToT + Prompt Refined (Yao et al., 2023) 7.91
Meta-Reasoner 7.68
The results are shown in Table 7. Meta-Reasoner achieves a coherency score of 7.68, which is competitive with specialized methods like ToT + Refine (7.91) and significantly outperforms basic approaches such as IO (6.19) and CoT (6.93). These results indicate that our framework effectively transfers reasoning capabilities to creative tasks where objectives are less precisely defined, thereby supporting the claim that our method is indeed domain-agnostic.
5.4.5. Stability of Dynamic Strategy Generation.
Dynamically expanding or refining the strategy set, as outlined in Section 4.3, enhances adaptability but poses risks of instability due to potential random strategy generation. To mitigate this, our method incorporates three stabilizing mechanisms:
(1) Initial Stable Foundation: We start with a curated set of verified strategies (e.g., Table 1), such as $g_{1}=$ ”Pause to clarify and disambiguate reasoning” and $g_{2}=$ ”Decompose task into sub-tasks,” ensuring a robust baseline. The LLM generates new strategies $g_{t}$ based on the current progress report $P_{t}$, constrained by contextual relevance, preventing arbitrary additions.
(2) Exploration-Exploitation Balance: The CMAB employs an $\epsilon$-greedy policy to balance exploration and exploitation, with the probability of selecting a new arm ($a_{t}\in G_{t}\setminus G_{t-1}$) governed by $\epsilon_{t}=\frac{1}{t}$. Suboptimal strategies are filtered out during inference based on their empirical reward $R_{t}(a_{t},P_{t})$. For instance, on the Game-of-24 task, the dynamic bandit achieves 89% accuracy, outperforming fixed-strategy sets (65%–72%), as shown in Table 6, demonstrating effective prioritization of viable strategies.
(3) Reward-Driven Feedback: The reward function $R:G_{t}\times P_{t}\to[0,1]$ (detailed in Section 4.3) provides real-time evaluation of strategy efficacy. Strategies yielding low rewards (e.g., $R_{t}<0.3$)) are deprioritized, ensuring rapid elimination of ineffective options. Figure 3 illustrates a monotonic increase in cumulative reward $\sum_{i=1}^{t}R_{i}$, reflecting stable reasoning as new strategies are integrated.
Collectively, these mechanisms ensure the dynamic CMAB remains stable by anchoring strategy generation in context, filtering via bandit optimization, and refining through reward feedback. Empirical results on the Game-of-24 and TheoremQA datasets (Table 6) confirm that the adaptive benefits of dynamic strategy generation outweigh instability risks, with performance gains of 17%–24% over static approaches.
6. Conclusion